# Why Go Full? Elevating Federated Learning Through Partial Network Updates

### v2

@@ Abstract
Federated learning is a distributed machine learning paradigm designed to protect user data privacy, which has been successfully implemented across various scenarios. In traditional federated learning, the entire parameter set of local models is updated and averaged in each training round. Although this full network update method maximizes knowledge acquisition and sharing for each model layer, it prevents the layers of the global model from cooperating effectively to complete the tasks of each client, a challenge we refer to as layer mismatch. This mismatch problem recurs after every parameter averaging, consequently slowing down model convergence and degrading overall performance. To address the layer mismatch issue, we introduce the FedPart method, which restricts model updates to either a single layer or a few layers during each communication round. Furthermore, to maintain the efficiency of knowledge acquisition and sharing, we develop several strategies to select trainable layers in each round, including sequential updating and multi-round cycle training. Through both theoretical analysis and experiments, our findings demonstrate that the FedPart method significantly surpasses conventional full network update strategies in terms of convergence speed and accuracy, while also reducing communication and computational overheads.
@@ 1 Introduction
Although this paradigm has been successful in many scenarios (Hard et al., 2018; Rieke et al., 2020), its convergence speed and ultimate performance are often lower than those of centralized schemes (McMahan et al., 2017; Zou et al., 2023), even when data across clients are independently and identically distributed (i.i.d.). This suggests that while full network updates and sharing enrich each model layer with more knowledge, they also introduce potential negative effects on final performance. To further investigate the underlying reason, we conduct an experiment to visualize the update step sizes during each iteration. Typically, in centralized learning, the update step sizes of the model show a downward trend, indicating that the model is gradually converging. However, in federated learning, as shown in Fig. 1a, the update step sizes significantly increase after each parameter averaging. This suggests that after averaging, the gradients calculated by subsequent layers become particularly large, indicating inadequate cooperation among layers within the global model, a phenomenon we term as layer mismatch. The cause of this issue is illustrated in Figure 2a. The middle section of the figure depicts the local models of each client, which have undergone sufficient local training. Within these local models, the layers cooperate effectively, demonstrating match. However, upon aggregating the parameters of each layer, the averaged layers may struggle to maintain this cooperation, resulting in mismatch. Such layer mismatch could lead to two problems: firstly, the final global model in federated learning may not converge to the optimal point of the global loss function, thereby negatively affecting performance. Secondly, the overall federated learning process may continually experiencing disruption due to mismatch by the server, substantially hampering training efficiency.
@@
To address the aforementioned problems, we propose FedPart, which employs partial network updates. Our main motivation is illustrated in Fig. 2b. In this toy example, we assume that only the $i$-th layer of the network is trainable in the current round. During local training of each client, the $i$-th layer can naturally align with other frozen parameters. These untrainable layers act as anchors, constraining the update direction of the $i$-th layer. This makes the averaged layers more likely to align with other layers. To validate this approach, we conduct experiments and visualize the results in Figure 1b. The curves clearly demonstrate that partial network updates significantly reduce the increase in update step sizes after averaging, thereby affirming their effectiveness in mitigating layer mismatch.
@@
However, as a trade-off, training and transmitting only a portion of the parameters at a time might limit the efficiency of knowledge learning and sharing. After thorough analysis, we identify that the solution to this problem lies in the strategy for selecting trainable parameters. Therefore, we carefully design a strategy for selecting trainable parameters in FedPart, drawing on two key principles. The first principle is sequential updating. We train the network layers sequentially, from shallow to deep, one layer at a time. This design is based on the observation that the shallower layers of a neural network typically converge to their final parameters faster than the deeper ones (Raghu et al., 2017). To maintain consistency with this inherent order, we apply a similar sequential strategy in layer selection. The second principle is the multi-round cycle training strategy. Our method emphasizes the importance of repeating the process of training from shallow layers to deep layers multiple times. During the original full network updates, shallow layers often learn low-level features, while deep layers learn high-level semantic features (Zeiler and Fergus, 2014; Erhan et al., 2009). To preserve this property, inspired by the idea of Block Coordinate Descent (BCD) (Poczos and Tibshirani, ), we propose the multi-round cycling training method to retain this characteristic to the greatest extent.
@@
In addition, FedPart also features computation and communication efficiency, which makes it highly suitable for edge computation scenarios (Wang et al., 2019a, b; Abreha et al., 2022). This is because FedPart only needs to train a part of the neural network at each training round, thereby significantly reducing the computational overhead of each client in each iteration. At the same time, since clients only need to upload and download the parts of the model that need updating, the amount of parameters to be transmitted is also greatly reduced.
@@
To validate the effectiveness of FedPart, we conduct explorations from both theoretical and experimental perspectives. Theoretically, we demonstrate that FedPart has a superior convergence rate under non-convex settings compared to FedAvg. Experimentally, we perform extensive evaluations on various datasets and model architectures. The results indicate that the FedPart method significantly improves convergence speed and final performance (e.g., an improvement of 24.8% on Tiny-ImageNet with ResNet-18), while also reducing both communication overhead (by 85%) and computational overhead (by 27%) simultaneously. Furthermore, our ablation experiments highlight the individual contributions of the proposed strategies in enhancing the overall performance of FedPart. We also include comprehensive visualization experiments to demonstrate the internal rationality of FedPart. In summary, the contributions of this paper are as follows:
@@
  • We observe the issue of layer mismatch in federated learning, caused by updating and aggregating all parameters in each training round. This phenomenon can potentially impact the model’s convergence speed and overall performance.
@@
  • We analyze the convergence rate of FedPart in a non-convex setting, demonstrating its theoretical advantages over full network updates.
  • We perform extensive experiments, showing that FedPart achieves significant improvements across multiple evaluation metrics compared to the full network update scheme. Additionally, ablation and visualization experiments help to further understand the rationality of FedPart.
@@ 2 Related Work
Currently, researches on partial parameter training or aggregation in federated learning has seen some applications, which can be broadly categorized into three types:
Train all parameters, aggregate partial parameters. Also known as personalized federated learning, this approach involves each client training all parameters but only aggregating some of them (Tan et al., 2022). For example, FedPer (Arivazhagan et al., 2019) and FedBN (Li et al., 2021b) personalizes classification and batch-normalization layers respectively, FedRoD (Chen and Chao, 2021) applies both global and local classifier heads. Other works may only upload a low-rank space of parameter matrices (Wang et al., 2023; Wu et al., 2024a). Although these methods achieve impressive results in data-heterogeneous scenarios, they usually exhibit performance degradation when datasets across clients are distributed in an i.i.d. manner. Additionally, these methods do not reduce computational overhead and, and the reduction in communication overhead is also relatively minimal since the personalized parts are usually small.
Train partial parameters, aggregate all parameters. This category refers to each client training a different part of the model, with a general update to the entire model during aggregation. For example, PVT (Yang et al., 2022) and FedPT (Sidahmed et al., 2021) strategically assigns specific model layers to each client, Federated Dropout (Caldas et al., 2018) and FedPMT (Wu et al., 2023) randomly assign a neurons in different layers to clients, HeteroFL (Diao et al., 2020) and FjORD (Horvath et al., 2021) deterministically decide a trainable subnetwork based on client computational power, FedRolex (Alam et al., 2022) further introduces a sliding window method, and CoCoFL (Pfeiffer et al., 2022) introduces a quantization technique for overhead reduction. The core goal of these methods is generally to reduce client overhead, especially focusing on dynamically utilizing the varying computational powers for clients. However, compared to full network updates, these methods often result in performance degradation and slower convergence speed.
Progressive Training. This approach starts with a small model and gradually increases its size until the entire network is trained (Rusu et al., 2016). This training paradigm has gain attention in the field of federated learning as its efficiency in reducing resource consumption (e.g., ProgFed (Wang et al., 2022) and ProFL (Wu et al., 2024b)). However, because these methods eventually train a full model, they are not able to solve the layer mismatch problem. Moreover, while they aim to reduce resource consumption, they often result in performance loss compared to full network training. To the best of our knowledge, our FedPart is the first to simultaneously boost convergence accuracy and efficiency.
@@ 3 Method
Generally speaking, FedPart is based on partial network updates, which trains and aggregates only few layers of the global network model for each training round. At the beginning of each training round that requires partial network update, the server first determines which layers need to be trained and sends this information to all clients. Subsequently, each client trains the corresponding layers, transmitting them to the server for aggregation, and the server broadcasts the averaged results to each client for next training round. We elaborate on two crucial elements in the subsequent subsections: partial network updates and the strategy for selecting trainable layers.
@@ 3.1 Partial Network Updates
The partial network updates involve training and aggregating only few layers of the global network model in each later communication rounds. Specifically, we partition the layers of global model into trainable ones and frozen ones. For each training iteration $t$ and client $i$, the optimization objective is: $\mathop{\arg\min}_{w_{i}^{t}}\mathbb{E}_{x\sim\mathbf{D_{i}}}[\mathcal{L}(x|\hat{w_{i}^{t}},\tilde{w}_{i}^{t})],$ where $\hat{w_{i}^{t}}$ and $\tilde{w}_{i}^{t}$ respectively denotes parameters of trainable and non-trainable layers, $\mathbf{D_{i}}$ represents the local data distribution of client $i$ and $\mathcal{L}(\cdot)$ refers to the loss function. To optimize this objective, we adopt the following gradient descent formula:
@@
Here, $w_{i}^{t}\triangleq\{\hat{w}_{i}^{t},\tilde{w}_{i}^{t}\}$ represents the total parameter set, $\gamma$ is the learning rate, $S_{i}^{t}$ is a binary mask that selectively enables updates only for trainable parameters and $\odot$ denotes element-wise product. After performing several local training iterations, the parameters of these selected layers are sent to the server and globally averaged at iteration $T$: $\bar{w}_{T}=\frac{1}{N}\sum_{i=1}^{N}w_{i}^{T},$ where $N$ represents the number of clients. In the formula mentioned above, for simplicity of formulation, we aggregate and calculate the gradient for all parameters. However, in practical implementation, we only update and transmit the trainable parts, thereby significantly reducing the computational and communication costs.
3.2 Strategy for Selecting Trainable Layers
Training only a subset of parameters in each round may restrict the efficiency of knowledge learning and sharing. Our main aim in designing the strategy for selecting trainable layers is to mitigate this limitation. As illustrated in Fig.3, following the initial full network updates, we train parameters layer by layer from the shallowest to the deepest. Subsequently, we cycle back to the shallowest layer and periodically repeat this process. This strategy encompasses the following two key principles:
@@ 3.2 Strategy for Selecting Trainable Layers
Sequential updating. This principle refers to training model layers in sequence, from shallow to deep layers one at a time. Our motivation is that the convergence of neural network exhibits a natural intrinsic order, with shallower layers typically converging earlier than deeper ones (Raghu et al., 2017). By updating partial network in accordance with this inherent training order, we can mirror the convergence process of full network updates, thus maintaining the training efficiency.
Multi-round cycle training. This principle refers to repeating the process of updating the neural network layers from shallow to deep multiple times. We illustrate our motivation with an example: in a fully trained neural network, shallow layers primarily focus on low-level semantic features (such as the edges in images), while deeper layers focus on higher-level semantic features (such as the main objects in images). However, during partial network updates, because the deeper layers are initially non-trainable, shallow layers are forced to learn complex high-level semantic features, which disrupts the original information hierarchy in the neural network. Through multi-round cycle training, we return to the shallow layers after training the deep layers. This strategy can reduce the burden on the shallow layers and approximate the final effect of full network updates.
@@ 3.3 Convergence Analysis for FedPart
To represent partial network updates, we add a binary matrix $S_{i}^{t}$ as a mask for each update process of each client. To keep consistency with the methods section, we assume that for each mask $S_{i}^{t}$, $\frac{1}{M}$ of the elements are 1, and the rest are 0. Before proving the convergence of our FedPart under non-convex conditions, we first propose three necessary assumptions:
@@
The first two assumptions are common in the literature, ensuring certain necessary characteristics of the loss function. The third assumption ensures that the specific choice of the mask matrix does not have too much impact on the final update variance, as long as the mask meets the requirements. A further discussion about the third assumption can be found in Appendix G.
Based on these assumptions, we can analyze the convergence of FedPart. In terms of the approximate convergence rate, we maintain consistency with related literature (Alistarh et al., 2017; Lian et al., 2017; Ghadimi and Lan, 2013), using the average magnitude of the expected gradient over iterations, and finally obtain the following theorem form:
Theorem: Under assumptions 1-3, with the total number of clients as $N$, and all parameters divided into $M$ groups of trainable parameters, the convergence rate of FedPart satisfies:
@@
where $\odot$ denotes element-wise product. A detailed proof of this theorem is provided in Appendix B. The results show that compared to the convergence rate of full network updates $O(\frac{1}{\sqrt{NT}})$ (Yu et al., 2019), FedPart’s convergence performance is significantly better. This advantage becomes more pronounced as the choice amount of partial parameters per instance is reduced, which aligns with our original intention to reduce layer mismatch. However, it should be noted that convergence analysis can only indicate the difficulty of converging to a stationary point, and cannot measure the model’s performance after convergence. Therefore, in practice, it is not advisable to indiscriminately reduce the amount of parameters trained each time.

### v3

@@ Abstract
Federated learning is a distributed machine learning paradigm designed to protect user data privacy, which has been successfully implemented across various scenarios. In traditional federated learning, the entire parameter set of local models is updated and averaged in each training round. Although this full network update method maximizes knowledge acquisition and sharing for each model layer, it prevents the layers of the global model from cooperating effectively to complete the tasks of each client, a challenge we refer to as layer mismatch. This mismatch problem recurs after every parameter averaging, consequently slowing down model convergence and degrading overall performance. To address the layer mismatch issue, we introduce the FedPart method, which restricts model updates to either a single layer or a few layers during each communication round. Furthermore, to maintain the efficiency of knowledge acquisition and sharing, we develop several strategies to select trainable layers in each round, including sequential updating and multi-round cycle training. Our theoretical analysis and experimental results show that the FedPart method significantly outperforms traditional full-network update strategies, achieving faster convergence, greater accuracy, and reduced communication and computational overhead.
@@ 1 Introduction
While this approach has proven effective in many applications (Hard et al., 2018; Rieke et al., 2020), its convergence speed and ultimate performance are often lower than those of centralized schemes (McMahan et al., 2017; Zou et al., 2023), even when data across clients are independently and identically distributed (i.i.d.). This suggests that while full network updates and sharing enrich each model layer with more knowledge, they may also introduce factors that negatively impact final performance. To further investigate the underlying reason, we conduct an experiment to visualize the update step sizes during each iteration. Typically, in centralized learning, the update step sizes of the model show a downward trend, indicating that the model is gradually converging. However, in federated learning, as shown in Fig. 1a, the update step sizes significantly increase after each parameter averaging. This suggests that after averaging, the gradients calculated by subsequent layers become notably large, indicating inadequate cooperation among layers within the global model, a phenomenon we term as layer mismatch. The cause of this issue is illustrated in Figure 2a. The middle section of the figure depicts the local models of each client, which have undergone sufficient local training. Within these local models, the layers cooperate effectively, demonstrating match. However, upon aggregating the parameters of each layer, the averaged layers may struggle to maintain this cooperation, resulting in mismatch. This layer mismatch can result in two key issues: first, it may prevent the global model in federated learning from converging to the optimal point of the global loss function, thereby compromising performance. Second, the persistent mismatch disrupts the federated learning process at the server level, significantly reducing training efficiency.
@@
To address the aforementioned problems, we propose FedPart, which employs partial network updates. Our main motivation is illustrated in Fig. 2b. In this toy example, we assume that only the $i$-th layer of the network is trainable in a given round. During local training on each client, this trainable layer can naturally align with the fixed parameters of other layers, which serve as anchors that constrain its update direction. This makes the averaged layers better align with other layers. To validate this approach, we conduct experiments and visualize the results in Figure 1b. The curves demonstrate that partial network updates significantly reduce the increase in update step sizes after averaging, confirming their role in alleviating layer mismatch.
@@
However, as a trade-off, training and transmitting only a portion of the parameters at a time might limit the efficiency of knowledge learning and sharing. Through thorough analysis, we identify a solution to this challenge, which involves carefully selecting trainable parameters. This solution is based on two key principles. The first principle is sequential updating. We train the network layers sequentially, from shallow to deep, one layer at a time. This design is based on the observation that the shallower layers of a neural network typically converge to their final parameters faster than the deeper ones (Raghu et al., 2017). To align with this natural order, we adopt a similar sequential strategy for layer selection. The second principle is the multi-round cycle training strategy. Our method emphasizes the importance of repeating the process of training from shallow layers to deep layers multiple times. During the original full network updates, shallow layers often learn low-level features, while deep layers learn high-level semantic features (Zeiler and Fergus, 2014; Erhan et al., 2009). To preserve this property, inspired by the idea of Block Coordinate Descent (BCD) (Poczos and Tibshirani, ), we propose the multi-round cycling training method to retain this characteristic to the greatest extent.
@@
In addition, FedPart offers improved computation and communication efficiency, making it highly suitable for edge computing scenarios (Wang et al., 2019a, b; Abreha et al., 2022). This is because FedPart only needs to train a part of the neural network at each training round, thereby significantly reducing the computational overhead of each client in each iteration. At the same time, since clients only need to upload and download the parts of the model that need updating, the amount of parameters to be transmitted is also greatly reduced.
@@
To validate the effectiveness of FedPart, we explore its performance from both theoretical and experimental perspectives. Theoretically, we demonstrate that FedPart has a superior convergence rate under non-convex settings compared to FedAvg. Experimentally, we perform extensive evaluations on various datasets and model architectures. The results indicate that the FedPart method significantly improves convergence speed and final performance (e.g., an improvement of 24.8% on Tiny-ImageNet with ResNet-18), while also reducing both communication overhead (by 85%) and computational overhead (by 27%) simultaneously. Furthermore, our ablation experiments demonstrate how each of the proposed strategies contributes to enhancing the overall performance of FedPart. We also conduct comprehensive visualization experiments to illustrate the underlying rationale of FedPart. In summary, the contributions of this paper are as follows:
@@
  • We identify the issue of layer mismatch in federated learning, which arises from updating and aggregating all parameters in each training round. This phenomenon can potentially impact the model’s convergence speed and overall performance.
@@
  • We theoretically analyze the convergence rate of FedPart in a non-convex setting, demonstrating its advantages over full network updates.
  • We perform extensive experiments, showing that FedPart achieves significant improvements across multiple evaluation metrics compared to the full network updates scheme. Additionally, ablation and visualization experiments enhance our understanding of the rationale behind FedPart.
@@ 2 Related Work
Current research on partial parameter training or aggregation in federated learning has led to various applications, broadly categorized into three types:
Train all parameters, aggregate partial parameters. Also known as personalized federated learning, this approach involves each client training all parameters but only aggregating some of them (Tan et al., 2022). For example, FedPer (Arivazhagan et al., 2019) and FedBN (Li et al., 2021b) personalizes classification and batch-normalization layers respectively, FedRoD (Chen and Chao, 2021) applies both global and local classifier heads. Other works may only upload a low-rank space of parameter matrices (Wang et al., 2023; Wu et al., 2024a). Although these methods achieve impressive results in data-heterogeneous scenarios, they usually exhibit performance degradation when datasets across clients are distributed in an i.i.d. manner. Moreover, they do not effectively reduce computational overhead, and any reduction in communication overhead is minimal, as the personalized components are typically small.
Train partial parameters, aggregate all parameters. This category refers to each client training a different part of the model, with a global update to the entire model during aggregation. For example, PVT (Yang et al., 2022) and FedPT (Sidahmed et al., 2021) strategically assigns specific model layers to each client, Federated Dropout (Caldas et al., 2018) and FedPMT (Wu et al., 2023) randomly assign a neurons in different layers to clients, HeteroFL (Diao et al., 2020) and FjORD (Horvath et al., 2021) deterministically decide a trainable subnetwork based on client computational power, FedRolex (Alam et al., 2022) further introduces a sliding window method, and CoCoFL (Pfeiffer et al., 2022) introduces a quantization technique for overhead reduction. The primary objective of these methods is to reduce client overhead by dynamically leveraging varying computational capacities among clients. However, compared to full network updates, these methods often result in performance degradation and slower convergence speed.
Progressive Training. This approach starts with a small model and gradually increases its size until the entire network is trained (Rusu et al., 2016). This training paradigm has gain attention in the field of federated learning as its efficiency in reducing resource consumption (e.g., ProgFed (Wang et al., 2022) and ProFL (Wu et al., 2024b)). However, because these methods eventually train a full model, they are not able to solve the layer mismatch problem. Moreover, while these methods aim to reduce resource consumption, they often lead to performance losses compared to full network training. To the best of our knowledge, our FedPart is the first approach to simultaneously enhance both convergence accuracy and efficiency.
@@ 3 Method
Generally speaking, FedPart is based on partial network updates, which trains and aggregates only a few layers of the global network model for each training round. At the beginning of each training round that requires partial network update, the server first determines which layers need to be trained and sends this information to all clients. Subsequently, each client trains the corresponding layers, transmitting them to the server for aggregation, and the server broadcasts the averaged results to each client for next training round. We elaborate on two key components of FedPart in the following subsections: partial network updates and the strategic selection of trainable layers.
@@ 3.1 Partial Network Updates
The partial network updates involve training and aggregating only a few layers of the global network model in each later communication rounds. Specifically, we partition the layers of global model into trainable ones and frozen ones. For each training iteration $t$ and client $i$, the optimization objective is: $\mathop{\arg\min}_{w_{i}^{t}}\mathbb{E}_{x\sim\mathbf{D_{i}}}[\mathcal{L}(x|\hat{w_{i}^{t}},\tilde{w}_{i}^{t})],$ where $\hat{w_{i}^{t}}$ and $\tilde{w}_{i}^{t}$ respectively denotes parameters of trainable and non-trainable layers, $w_{i}^{t}\triangleq\{\hat{w}_{i}^{t},\tilde{w}_{i}^{t}\}$ represents the total parameter set, $\mathbf{D_{i}}$ represents the local data distribution of client $i$ and $\mathcal{L}(\cdot)$ refers to the loss function. To optimize this objective, we adopt the following gradient descent formula:
@@
Here, $\gamma$ is the learning rate, $S_{i}^{t}$ is a binary mask that selectively enables updates only for trainable parameters and $\odot$ denotes element-wise product. After performing several local training iterations, the parameters of these selected layers are sent to the server and globally averaged at iteration $t=T$: $\bar{w}_{T}=\frac{1}{N}\sum_{i=1}^{N}w_{i}^{T},$ where $N$ represents the number of clients. For the sake of simplicity in formulation, the above equation aggregates and calculates the gradient for all parameters. However, in practical implementation, we only update and transmit the trainable components, significantly reducing both computational and communication costs.
3.2 Selecting Trainable Layers
Although training only a subset of parameters can largely mitigate the layer-mismatch issue, it may limit the efficiency of knowledge learning and sharing. Therefore, we propose to carefully select trainable layers to address this limitation. As illustrated in Fig.3, following the initial full network updates, we train parameters layer by layer from the shallowest to the deepest. Subsequently, we cycle back to the shallowest layer and periodically repeat this process. Generally, this strategy is driven by two key principles:
Sequential updating. This principle refers to training model layers in sequence, from shallow to deep layers one at a time. Our motivation is based on the fact that the convergence of neural networks follows a natural intrinsic order, with shallower layers typically converging earlier than deeper ones (Raghu et al., 2017). By partially updating network in accordance with this inherent training order, we can largely replicate the convergence process of full network updates while preserving training efficiency simultaneously.
Multi-round cycle training. This principle refers to repeating the process of updating the neural network layers from shallow to deep multiple times. To illustrate our motivation, consider a fully trained neural network: shallow layers primarily focus on low-level semantic features (such as the edges in images), while deeper layers focus on higher-level semantic features (such as the main objects in images). However, during partial network updates, because the deeper layers are initially non-trainable, shallow layers are forced to learn complex high-level semantic features, which disrupts the original information hierarchy in the neural network. Through multi-round cycle training, we return to the shallow layers after training the deep layers. This strategy can reduce the burden on the shallow layers and helps approximate the final results of full network updates.
@@ 3.3 Convergence Analysis for FedPart
To represent partial network updates, we add a binary matrix $S_{i}^{t}$ as a mask for each update process of each client. To keep consistency with the methods section, we assume that for each mask $S_{i}^{t}$, $\frac{1}{M}$ of the elements are set to 1, and the rest are set to 0. Before proving the convergence of our FedPart under non-convex conditions, we first introduce three assumptions:
@@
The first two assumptions are common in the literature, ensuring certain necessary characteristics of the loss function. The third assumption ensures that the specific choice of the mask matrix does not have too much impact on the final update variance, provided that the mask meets the required criteria. A further discussion about the third assumption can be found in Appendix G.
Based on these assumptions, we can analyze the convergence of FedPart. In terms of the approximate convergence rate, we align with related literature (Alistarh et al., 2017; Lian et al., 2017; Ghadimi and Lan, 2013), using the average magnitude of the expected gradient over iterations, and finally obtain the following theorem:
Theorem 1: Under assumptions 1-3, with the total number of clients as $N$, and all parameters divided into $M$ groups of trainable parameters, the convergence rate of FedPart satisfies:
@@
where $\odot$ denotes element-wise product. A detailed proof of this theorem is provided in Appendix B. The results show that compared to the convergence rate of full network updates $O(\frac{1}{\sqrt{NT}})$ (Yu et al., 2019), FedPart demonstrates significantly better convergence performance. This advantage becomes more pronounced as the choice amount of partial parameters per instance is reduced, which aligns with our original intention to reduce layer mismatch. However, it should be noted that convergence analysis can only indicate the difficulty of converging to a stationary point, and cannot measure the model’s performance after convergence. Therefore, it is not advisable to arbitrarily reduce the number of parameters trained in each iteration.